# Bar properties as a function of wavelength in Illustris TNG50: analysis of mock images

### v1

@@ 4.2 Mass maps are similar to infrared filters
Taking this aspect into consideration, one of the most important results of our paper is the comparison between measurements of bar properties in stellar mass maps and in observational filters, the latter resulting from radiative transfer mocks. We emphasize that in the context of cosmological simulations, comparing the filters in relation to the mass distribution has particularly useful results, in some cases more useful than comparing the filters with each other. This result is clearly illustrated by the solid black line in Fig. 4, which shows that the mass map presents ellipticity and length values characteristic of infrared filters, specifically aligning with those from the Spitzer IRAC/$3.6\,\upmu\text{m}$ channel. It is understandable that the mass maps resemble our infrared filter, as this wavelength is notably less affected by extinction effects and more accurately represents the large population of low-mass stars that dominate the galactic disc (Eskridge et al., 2000). These stars, although individually less massive, collectively dominate the galactic disc due to the distributions prescribed by the Initial Mass Function (IMF), as detailed in Chabrier (2003). Unlike blue filters, which prominently display young stellar regions with active star formation, the infrared view focuses on the broader, underlying mass distribution, where such actively star-forming regions are less emphasized.
@@ 5 Summary and Conclusion
We created realistic mock images of 50 barred galaxies from Illustris TNG50, using filters from infrared to UV, and measured the bar properties of this sample using ellipse fitting. The main results of this paper can be summarized as follows:
@@
    For bar length, the mean difference across filters is inconclusive when regarding the entire sample. This is due to passive galaxies with short bars, that drive the inverted trend. When only star-forming galaxies are considered, then we find a statistically significant mean difference in bar length across filters. Bar length, in this star-forming subsample, increases by 10 per cent and 17 per cent from the $3.6\,\upmu\text{m}$ to the $g$ and $J0378$ bands, respectively. This behaviour aligns with the findings of Menéndez-Delmestre et al. (2024).
@@ Appendix B Mock Images of the Galaxy Sample
Figure 19: Mass maps and mock images of our galaxy sample in different wavebands: Spitzer IRAC/$3.6\,\upmu\text{m}$, SDSS i, SDSS r, SDSS g, SPLUS J0378, Galex NUV, and Galex FUV. All the figures were generated with a size of 60 $\times$ 60 kpc in each frame and with a face-on orientation.

### v2

@@ 4.2 Mass maps are similar to infrared filters
[…]
This similarity is to be expected, as infrared wavelengths are less affected by extinction and more accurately trace the large population of low-mass stars that dominate the galactic disc (Eskridge et al., 2000). These stars, although individually less massive, collectively dominate the galactic disc due to the distributions prescribed by the Initial Mass Function (IMF), as detailed in Chabrier (2003). Unlike blue filters, which prominently display young stellar regions with active star formation, the infrared view focuses on the broader, underlying mass distribution, where such actively star-forming regions are less emphasized.
@@ 5 Summary and Conclusion
We created realistic mock images of 50 barred galaxies from TNG50, using filters from infrared to UV, and measured the bar properties of this sample using ellipse fitting. The main results of this paper can be summarized as follows:
@@
    For bar length, the mean difference across filters is inconclusive when regarding the entire sample. This is due to passive galaxies showing shorter bars in bluer filters (e.g. $J0378$ and $g$), which drives the inverted trend, with an average slope approximately half that of the star-forming group, meaning the bar-length difference across filters is smaller. When only star-forming galaxies are considered, then we find a statistically significant mean difference in bar length across filters. Bar length, in this star-forming subsample, increases by 10 per cent and 17 per cent from the $3.6\,\upmu\text{m}$ to the $g$ and $J0378$ bands, respectively, and this behaviour aligns with the findings of Menéndez-Delmestre et al. (2024).
@@ Appendix B Mock Images of the Galaxy Sample
[…]